An administrator at a college claims that the mean SAT Mathematics score of incoming students is 520. You find that in a random sample of 45 incoming students, the mean SAT Mathematics score is 511 with a standard deviation of 48.65. Assume the population of scores are normally distributed.  

Suppose you perform a hypothesis test to determine whether the mean SAT Mathematics score of incoming students is less than 520.

What is the P-value for this hypothesis test? Round the value to two decimal places.
Answer #1

Solution :

The null and alternative hypothesis are

H0 : \mu = 520   ........... Null hypothesis

Ha : \mu < 520   ........... Alternative hypothesis

Here,  n = 45, \bar x = 511, s = 48.65

The test statistic t is,

t =  T-]/[s/n

= [511 - 520]/[48.65 /\sqrt{}45]

= -1.24

The value of the test​ statistic t = -1.24

Now ,

d.f. = n - 1 = 45 - 1 = 44

< sign in Ha indicates that the test is ONE TAILED.

t = -1.24

So , using calculator ,

p value = 0.1108

Answer : The P-value for this hypothesis test is 0.11